# Uncovering Bias Paths with LLM-guided Causal Discovery: An Active Learning and Dynamic Scoring Approach

### v1

@@ 10 Limitations
Simplified Causal Modeling in Synthetic Data.
Our synthetic graph, based on the UCI Adult dataset, models income as a terminal node to isolate fairness-related pathways. This design aids in evaluating methods’ ability to detect direct and indirect influences on income. However, it omits downstream effects that income may have in reality—such as on health, wealth, or opportunity, thus excluding feedback dynamics and longitudinal consequences. Future work could extend this framework to capture post-outcome influences and assess fairness in more dynamic causal settings.
Hyperparameter Sensitivity and Computational Constraints.
Although effective, our method requires careful tuning of multiple hyperparameters (e.g., scoring weights, thresholds, LLM temperature), demanding domain expertise. LLM-based querying also incurs substantial computational cost, especially for large networks. Token limits restrict the ability to encode full metadata, particularly in complex graphs like Neuropathic, leading to incomplete or inconsistent responses. Specialized variable names may be misinterpreted, especially at higher temperatures, introducing semantic errors and prediction noise. These factors collectively pose scalability and usability challenges, particularly for resource-constrained settings or high-dimensional domains.
Reproducibility Challenges in LLM-Based Causal Discovery.
Discrepancies between our results and those reported by Jiralerspong et al. (2024) highlight broader reproducibility concerns in LLM-based CD. LLMs are inherently stochastic and sensitive to prompt design, versioning, query order, and API behavior, factors rarely standardized or reported. Additionally, hidden dependencies on tokenization, temperature, or latent state carryover can affect outcomes, making benchmarking difficult. These challenges underscore the need for greater transparency, reproducible prompt templates, and controlled evaluation protocols in future work.
Risk of Social Bias in LLM Responses.
LLMs trained on web-scale data may encode social biases, leading to causal inferences that reflect stereotypes rather than ground-truth mechanisms. This risks semantic hallucinations and biased edge assignments, particularly in fairness-critical applications.
Domain Dependency and Metadata Limitations.
A further limitation lies in the domain dependency of our approach. Because our framework relies on variable metadata to guide LLM-based querying, it is best suited to domains with well-defined, semantically meaningful feature names (e.g., healthcare, census, education). In settings where metadata is sparse, ambiguous, or domain-specific (e.g., sensor networks or high-dimensional genomics), LLM-based judgments may degrade or hallucinate structure (E.g. Neuropathic network). Future work could explore hybrid prompting strategies or domain adaptation techniques to extend semantic CD to less interpretable or evolving data environments.
@@ 11 Conclusion
In this work, we propose a fairness-driven CD framework that integrates LLMs with AL and dynamic scoring. Our method outperforms baselines in recovering accurate causal structures across various benchmark networks, and demonstrates more conservative behavior in identifying fairness-relevant pathways, reducing false positives in bias attribution. By combining statistical signals with semantic priors, and prioritizing high-informative queries, the proposed framework advances the reliability and fairness-awareness of LLM-based CD.
To support robust evaluation, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding fairness-relevant causal structure, confounding, and noise. This benchmark provides a controlled yet realistic testbed for future fairness-oriented CD research.
Our hyperparameter analysis revealed that exploration depth (max iterations) strongly influences performance, while the importance of statistical versus LLM-derived signals varies with graph scale.
While our approach improves recovery of fairness-relevant pathways, our framework does not directly translate into end-to-end fairness guarantees in downstream tasks. Bridging the gap between discovery and actionable fairness interventions remains an important area for future integration. Other future directions include expanding to temporal graphs, refining prompt engineering for domain-specific metadata, and building reproducible pipelines for benchmarking LLM-guided causal inference.
@@
Appendix A Appendix: Synthetic Data Generation Rules
This section provides more detail on the Adult-based synthetic data general rules, formulas, and process.
@@ A.1 Data Generation Process
[…]
@@ A.2 Latent Confounding via Unobserved Variable $U$
To reflect unmeasured societal influences (e.g., privilege, institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$, that affects intermediate variables such as years of education, capital gain, and hours of work per week, thereby indirectly influencing income.
@@
[…]
Appendix B Appendix: Experimental Details
@@
We perform 200 trials per dataset, optimizing for F1 score via Bayesian search. The tuned hyperparameter space includes: query prioritization weights selected from $[(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)]$, score threshold values between 0.25 and 0.35, temperature values between 0.25 and 0.6, and a maximum iteration range of 40 to 100. We report the average across seeds for all performance metrics used in evaluation.
@@
Appendix C Appendix: Parameter Impact Analysis on Proposed Method For Child and Neuropathic Networks
@@
Appendix D Appendix: Fairness and Path Analyses Results
@@
The graphs generated using the proposed method, PC, GES, DAGMA, Notears, LLM Pairwise, and LLM BFS methods are shown in Figures 10 to 16.
@@
Appendix E Appendix: Complexity Analyses

### v2

@@ 9 Limitations
Simplified Synthetic Causal Structure.
Our synthetic graph models income as a terminal node to isolate fairness pathways, aiding evaluative clarity. However, this excludes realistic downstream effects (e.g., on health, opportunity), omitting feedback or temporal dynamics. Future work should extend this to longitudinal or post-outcome settings.
Hyperparameter Sensitivity and Compute Costs.
Our method requires tuning multiple hyperparameters (e.g., score weights, temperature), which can be resource-intensive. LLM querying is computationally costly, particularly for large graphs like Neuropathic. Token limits restrict the ability to encode full metadata, particularly in complex graphs like Neuropathic, leading to incomplete or inconsistent responses. Specialized variable names may be misinterpreted, especially at higher temperatures, introducing semantic errors and prediction noise. These factors collectively pose scalability and usability challenges, particularly for resource-constrained settings.
Reproducibility of LLM-based CD.
While our method is LLM-agnostic, we observe discrepancies with results from Jiralerspong et al. (2024), highlighting broader reproducibility challenges. LLMs are stochastic and sensitive to prompt design, versioning, query order, and API behavior, factors rarely standardized or reported. Without standardized prompting protocols or access to latent states, consistent benchmarking remains difficult.
Risk of Social Bias in LLMs.
LLMs trained on web-scale data may encode social biases, leading to causal inferences that reflect stereotypes rather than ground-truth mechanisms. This risks semantic hallucinations and biased edge assignments, particularly in fairness-critical applications. This risk is addressed through statistical weighting and confidence-based filtering, as described in Section 3.
Domain and Metadata Dependence.
Our framework depends on interpretable variable metadata to guide LLM queries. It performs best in domains like healthcare or census data. In sparse, technical, or ambiguous domains (e.g., genomics, sensor data), performance degrades. While LLM priors scale efficiently, human-in-the-loop priors could be incorporated for low-interpretability domains. Future work could explore hybrid prompting strategies or domain adaptation techniques to extend semantic CD to less interpretable or evolving data environments.
@@ 10 Conclusion
We propose a fairness-driven CD framework that integrates LLMs with AL. Across benchmarks including synthetic and semi-synthetic real-world networks, our method outperforms prior baselines in structural accuracy and reduces false positives in fairness-critical path recovery. By combining statistical dependencies with LLM-based semantic priors and prioritizing informative queries, our framework improves both reliability and fairness-awareness in CD.
To support reproducible evaluation, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding structural bias, latent confounding, and noise. This provides a realistic yet controlled testbed with ground-truth graphs for consistent comparison across CD methods.
Our analysis shows that exploration budget (e.g., number of iterations) significantly affects performance, with the relative importance of LLM-derived versus statistical signals shifting with graph size and complexity. While our method improves fairness-path recovery, it does not offer end-to-end guarantees for downstream fairness outcomes. Instead, it uncovers explainable causal pathways, particularly those involving sensitive attributes intended to guide audits, impact assessments, or stakeholder review.
This positions our framework as an improvement in diagnostic tools for fairness evaluations. Use cases include hiring audits, policy design, and regulatory compliance reporting. By recovering interpretable causal paths, our method offers a step toward tools that support fairness-focused exploration, potentially informing future integration into dashboards or algorithmic auditing pipelines. We plan to extend the framework to dynamic graphs and release open-source pipelines for scalable, reproducible fairness audits.
@@
Acknowledgments
This study was funded by NSF #2047296 and OpenAI’s Researcher Access Program.
Appendix A Synthetic Data Generation
This section provides more detail on the Adult-based synthetic data general rules, formulas, and process.The graph was constructed based on causal assumptions documented in prior fairness research (Kilbertus et al. 2017; Nabi and Shpitser 2018). Variable relationships were chosen to reflect widely accepted socioeconomic patterns and to ensure transparency and reproducibility, minimizing author-imposed bias.
@@ A.1 Data Generation Process
capital gain Numerical occupation $\mathcal{N}(5000+1500\cdot U,\ 2000)$ if occupation==1; else 1000 + 500 $\cdot$ U or 0
capital loss Numerical occupation $\mathcal{N}(1000,\ 500)$ if occupation==1; else 300 or 0
work hours per week Numerical occupation, sex $\mathcal{N}(45+2\cdot U,\ 5)$ if sex==0; else 40 + U or 35
final weight Numerical None $\mathcal{N}(180000,\ 20000)$
@@ A.2 Latent Confounding via Unobserved Variable $U$
To reflect unmeasured societal influences (e.g., privilege, institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$, that affects intermediate variables such as years of education, capital gain, and hours of work per week, thereby indirectly influencing income. This reflects latent societal advantages that are not captured in observed features.
@@
The latent confounder is injected via additive linear terms into structural equations following standard assumptions in structural causal models (SCMs). For the binary outcome income, we use a nonlinear threshold function that combines observable and latent contributions. This setup reflects bias amplification near decision boundaries and is consistent with prior fairness simulation frameworks.
@@
$\text{Income}=\mathbb{I}\Big[$ $\;\text{capital\_gain}+500\cdot U>4000$
$\land\;\text{education\_num}+0.5\cdot U\geq 14$
$\land\;\text{hours\_per\_week}+U>40\Big]$
A.3 Structured Noise Augmentation
Additive Noise for Continuous Variables.
For each continuous variable X, noise is sampled as:
Symmetric Label Noise in Categorical Variables
@@ A.3 Structured Noise Augmentation
This means with probability $\gamma$, the label is flipped uniformly to one of the K1K-1 incorrect categories. And with probability $1-\gamma$, the label remains correct. This uniform corruption model known as symmetric label noise, is widely used in label-noise robust learning (Natarajan et al. 2013; Patrini et al. 2017).
Controlled Replication Across Seeds
To evaluate generalization and fairness robustness, we generate four datasets using different random seeds (42, 101, 222, 333), each with the same underlying graph but distinct noise realizations. Noise is added post-graph generation, preserving causal structure while introducing rea)listic variability. This mirrors standard practices in fairness benchmarking and robustness testing for CD.
Appendix B Experimental Details
@@
We perform 200 trials per dataset, optimizing for F1 score via Bayesian search. The tuned hyperparameter space includes: query prioritization weights selected from [(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)][(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)], score threshold values between 0.25 and 0.35, temperature values between 0.25 and 0.6, and a maximum iteration range of 40 to 100. We report the average across seeds for all performance metrics used in evaluation.
@@
An example of a full prompt for this experiment is listed in Listing 2.
Listing 2 LLM Query Example - Syn-Adult Dataset
1System: You are an expert in socioeconomic and demographic research.
2User: You are a helpful assistant to an expert in socioeconomics, with a focus on income prediction and labor statistics.
3
4Metadata:
5Variable A: education: highest level of education achieved (e.g., Bachelors, Masters)
6Variable B: income: income level (whether the individuals income exceeds \\$50K/year) ...
7
8Query:
9Does education cause income?
10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
@@
An example of a full prompt for this experiment is listed in Listing 3.
Listing 3 LLM Query Example - CHILD Dataset
1System: You are an expert on childrens diseases.
2User: You are a helpful assistant to a childrens disease expert.
3
4Metadata:
5Variable A: BirthAsphyxia: lack of oxygen to the blood during the infants birth
6Variable B: Disease: infant methemoglobinemia
7
8Query:
9Does BirthAsphyxia cause Disease?
10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
@@
An example of a full prompt for this experiment is listed in Listing 4.
Listing 4 LLM Query Example - Neuropathic Dataset
1System: You are an expert on neuropathic pain diagnosis.
2User: You are a helpful assistant to a neuropathic pain diagnosis expert.
3
4Metadata:
5Variable A: L C6 Radikulopati: left C6 radiculopathy (nerve pain originating at cervical vertebra C6)
6Variable B: L Handbesvar: left hand discomfort
7
8Query:
9Does L C6 Radikulopati cause L Handbesvar?
10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
Appendix C Parameter Impact Analysis on Proposed Method For Child and Neuropathic Networks
@@
Appendix D Fairness and Path Analyses Results
@@
The graphs generated using the proposed method, PC, GES, DAGMA, Notears, LLM Pairwise, and LLM BFS methods are shown in Figures 9 to 15.
@@
Appendix E Complexity Analyses